OpenStudy (anonymous):

Write each expression as a power function. g(x)=3.7√3:x

OpenStudy (anonymous):

is this \[g(x)=3.7\sqrt[3]{x}\]?

OpenStudy (anonymous):

Yes thanks for helping. How do I do it?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[3.7timesx^(1/3)\]

OpenStudy (anonymous):

\[3.7\times x^{\frac{1}{3}}\] yes

OpenStudy (anonymous):

the the 1/3 is an exponent